# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

### v2

@@ Abstract
Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals. Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differential value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation. While CTRL has shown promise, its applications have been largely limited to the single-agent domain. This limitation stems from two key challenges: (i) conventional solution methods for HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with HJB-based learning approaches, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training. In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale. To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines $\nabla_{x}V$ along trajectories. This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning. We evaluate our method using continuous‑time variants of standard benchmarks, including multi‑agent particle environment (MPE) and multi‑agent MuJoCo. Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex multi-agent dynamics.
@@ 1 Introduction
RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation [5], strategy games [39], wireless communications [12, 44] , and traffic coordination [16]. Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval. However, discrete‑time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading [33] and autonomous driving [21]). Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues [41, 34]. To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time [9, 32]. However, existing works only focus on single‑agent settings and have not yet been widely explored for multi-agent scenarios. In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies. Within this context, our paper investigates a novel approach to solving CT-MARL problems. To better understand the limitations of DTRL and advantages of CTRL for multi-agent scenarios, we present a didactic case as shown in Fig. 1. In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2. In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
@@
Figure 1: The performance of CT-MARL and DT-MARL is compared on a continuous-time, two-agent coupled oscillator task. DT-MARL suffers from significant bias and error when applied in the continuous domain. In contrast, CT-MARL yields smoother actions, higher rewards, and more accurate value approximations, closely aligning with the analytical LQR ground truth.
@@
Unlike Bellman operator-based DTRL [2], CTRL leverages HJB PDEs to compute differential value functions [27, 34]. However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems [3], especially where the computational complexity grows exponentially with the state dimension in multi-agent systems. PINNs have emerged as a powerful tool to circumvent CoD [45], and offer convergence guarantees for problems with smooth solutions [35, 17]. To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging. The inherent non-stationarity across agents makes value learning unstable [49], and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients [51]. These inaccuracies propagate into the value function, undermining policy quality and stability. To address this limitation, we propose a novel learning approach that combines PINN and VGI information, which is used to optimize the value learning. The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories. This integration yields significantly more accurate value‑gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
@@
Our work makes the following contributions. (1) We leverage PINNs to approximate differential value functions and apply them to solve continuous‑time multi‑agent reinforcement learning problems, which have rarely been explored by previous studies. (2) We introduce a novel value‑gradient iteration term that dynamically refines the approximations of value gradients during training. This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning. (3) We create continuous‑time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo. The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value‑gradient learning in high‑dimensional multi‑agent systems.
@@ 2.1 Continuous-Time Reinforcement Learning
CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings [18, 23, 40, 47, 46, 38, 4, 11]. These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation. For instance, [4] proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy. Similarly, [11] designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees. [23] designs two policy iteration algorithms that compute comprehensive solutions to HJB equations. In contrast, [18] introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization. [47] develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics. [34] utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings. [48] integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization. CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings. [26] solves multi-agent pathfinding problem using fuzzy Q-learning, while [19] proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems. Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
@@ 2.2 Solving HJB Equations via PINNs
In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations [7], which are the first-order nonlinear parabolic PDEs. However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD [30, 29]. Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth [45]. PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual [14, 15], PDE residual [1, 51], and supervised data derived from numerical solvers [28, 50]. Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments [27]. However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
@@ 3.1 Problem Formulation
In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework [8, 48, 23], we formulate the continuous-time multi-agent problem as a tuple:
$\mathcal{M}=\bigl\langle\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,g,\{t_{k}\}_{k\geq 0},T,\rho\bigr\rangle.$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state space and $\mathcal{U}=\mathcal{U}_{1}\times\dots\times\mathcal{U}_{N}\subseteq\mathbb{R}^{m}$ represents the joint action space of $N$ agents. The global state and control input are represented by $x\in\mathcal{X}$ and $u\in\mathcal{U}$. Agent interactions occur over a fixed time horizon [0, T]. The multi-agent system evolves according to time-invariant nonlinear dynamics defined by $\dot{x}=f(x,u)$, where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is the global dynamics function. We define $\pi:\mathcal{X}\times[0,T]\rightarrow\mathcal{U}$ as the decentralized joint policy $\pi=(\pi_{1},\dots,\pi_{N})$. All agents share a global reward $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ and a global terminal cost $g:\mathcal{X}\rightarrow\mathbb{R}$. $\rho\in(0,1]$ is the discount factor or time-horizon scaling parameter. Unlike standard formulations that assume a fixed time step, we consider a strictly increasing sequence of decision times $\{t_{k}\}_{k\geq 0}$ with variable gaps $\tau_{k}=t_{k+1}-t_{k}>0$. In this paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f$ is Lipschitz continuous; $r$ and $g$ is Lipschitz continuous and bounded.
@@ 3.2 HJB and Policy Learning
For CT-MARL problems, we adopt the HJB theory to solve continuous-time problems. In this subsection, we explain how the HJB equation is leveraged to solve the CT-MARL problems. Specifically, we define a value network $V_{\theta}$ parametrized by a set of weights and biases $\theta$, and describe how the global value function $V_{\theta}$ is trained and how each agent’s policy $\pi_{\phi_{i}}$ is updated, such that the overall procedure serves as a continuous-time analogue of actor–critic policy iteration.
3.2.1 Critic Learning with HJB.
First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq. (2), is the optimal solution to HJB PDEs [7] in the context of cooperative multi-agent settings.
@@
For all $(x,t)\in\mathcal{X}\times[0,T]$, the value function $V(x,t)$ is the optimal solution to satisfy the HJB PDEs and boundary conditions:
$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f\bigl(x,u^{*}\bigr)+r\bigl(x,u^{*}\bigr)=0,$ (3)
where optimal control input $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, and boundary condition is $V(x,T)=g(x)$. The Hamiltonian $\mathcal{H}$ is defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f\bigl(x,u\bigr)+r\bigl(x,u\bigr)$.
@@
To approximate differential value functions, we aim to solve the HJ PDEs given in Eq. (3). However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 [6]. To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions. Specifically, we define HJB PDE residual as
$\mathcal{R}_{\theta}(x,t)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u).$ (4)
and minimize both residual loss and boundary condition loss towards zero during model refinement.
$\mathcal{L}_{\text{res}}=\bigl\|\mathcal{R}_{\theta}(x,t)\bigr\|_{1},\quad\mathcal{L}_{\text{term}}=\bigl\|V_{\theta}(x,T)-g(x)\bigr\|_{2}^{2}.$ (5)

### v3

@@ Abstract
Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals. Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differentiable value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation. While CTRL has shown promise, its applications have been largely limited to the single-agent domain. This limitation stems from two key challenges: (i) conventional methods for solving HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with learning-based approaches to alleviate the CoD, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training. In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale. To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines value gradients along trajectories. This improves gradient accuracy, in turn yielding more precise value approximations and stronger policy learning. We evaluate our method using continuous‑time variants of standard benchmarks, including multi‑agent particle environment (MPE) and multi‑agent MuJoCo. Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex cooperative multi-agent dynamics. Code is available at this link
@@ 1 Introduction
RL has achieved remarkable success in a range of discrete-time single- and multi-agent interaction tasks, including robotic manipulation (Brunke et al., 2022), strategy games (Vinyals et al., 2019), wireless communications (Feriani and Hossain, 2021; Wang et al., 2023), and traffic coordination (Haydari and Yılmaz, 2020). However, many real-world domains are inherently continuous-time and operate at high or irregular decision frequencies, such as continuous control (Talbot et al., 2025), autonomous driving (Brechtel et al., 2014), and market trading (Shavandi and Khedmati, 2022). Despite this, most existing RL methods formulate decision-making in discrete-time, where Bellman updates are computed at a fixed time interval. Such discrete-time RL (DTRL) approximates a continuous-time process by imposing a fixed discretization step, which introduces two inherent limitations. First, when the timestep is coarse, the resulting controller becomes non-smooth and leads to suboptimal or unstable behavior (Doya, 2000). Second, when the timestep is fine, the number of states and iteration steps become large, which enlarges not only memory storage but also many learning trials (Doya, 2000). In addition, as the time step $\Delta t$ approaches to 0, the Bellman operator can become ill-conditioned: the temporal-difference objective functions may be dominated by approximation noise, leading the Bellman updates to become unstable and have a large variance (Wang et al., 2020; Shilova et al., 2024). These limitations indicate that performance may critically depend on the chosen discretization method, thereby motivating continuous-time RL (CTRL) methods that avoid timestep-discretization and directly learn value functions in continuous time (Doya, 2000; Rubanova et al., 2019).
@@
Figure 1: The performance of our CT-MARL and DT-MARL is compared on a continuous-time, two-agent coupled oscillator task. In the discrete-time setting, DT-MARL trained with MADDPG can achieve near-optimal performance. However, when transferred to the continuous-time domain, MADDPG suffers from significant bias and error, resulting in poor approximations. In contrast, CT-MARL yields smoother actions, higher rewards, and more accurate value approximations, closely aligning with the analytical LQR ground truth.
@@
Unlike Bellman operator-based DTRL (Bellman, 1966), CTRL leverages HJB Partial Differential Equation (PDEs) to compute differential value functions (Mukherjee and Liu, ; Shilova et al., 2024). However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or level set method (Osher et al., 2004)) suffers from CoD in high-dimensional dynamical systems (Bellman et al., 1965), especially where the computational complexity grows exponentially with the state dimension in multi-agent systems. PINNs have emerged as a powerful tool to circumvent CoD (Weinan et al., 2021), and offer convergence guarantees for problems with smooth solutions (Shin et al., 2020; Ito et al., 2021). To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
Yet existing CTRL methods focus almost exclusively on single-agent settings and do not extend naturally to MARL. In multi-agent domains, each agent must interact with the environment while simultaneously adapting to other learning agents, introducing severe non-stationarity, making value learning substantially harder (Yu et al., 2022). Even with PINNs, accurately learning value functions under the centralized training decentralized execution (CTDE) framework remains challenging: PDE-based residual losses alone often produce biased or inaccurate value gradients, where coupled dynamics amplify approximation errors (Zhang et al., 2024b). These inaccuracies propagate through the PDE residuals, degrade value learning, and ultimately deteriorate policy optimization. To address this limitation, we propose a novel learning approach that combines PINN and VGI to optimize the value learning. The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories. To better understand the limitations of DTRL and advantages of our CTRL algorithms for multi-agent scenarios, we present a didactic case as shown in Fig. 1. In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2. In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
@@
Our work makes the following contributions. (1) We leverage PINNs to approximate differential value functions and apply them to solve cooperative continuous‑time multi‑agent reinforcement learning problems within deterministic systems, which have rarely been explored by previous studies. (2) We introduce a novel value‑gradient iteration term that dynamically refines the approximations of value gradients during training. This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning. (3) We create continuous‑time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo. The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value‑gradient learning in high‑dimensional multi‑agent systems.
@@ 2.1 Continuous-Time Reinforcement Learning
CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings. These studies aim to optimize policies in continuous-time domains, avoiding time discretization and yielding more accurate control actions in robotics and navigation. For instance, Bian and Jiang (2021) proposes a continuous-time value iteration algorithm for solving HJB equations without the need for a stabilizing initial policy, while Faradonbeh and Faradonbeh (2023) introduces an HJB-based actor-critic network with theoretical guarantees for the infinite-horizon case. Similarly, Lee and Sutton (2021) develops two policy iteration algorithms that compute comprehensive solutions to HJB equations. In contrast, Jia and Zhou (2022a) introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization. Building on this line, Jia and Zhou (2022b; 2023) uses rigorous algorithms and strong theoretical foundations to study single-agent CTRL with stochastic dynamics. Other approaches include Yang et al. (2021), which develops a robust actor–critic framework for nonlinear systems with unmodeled dynamics; Shilova et al. (2024), which leverages PINNs with $\epsilon$-scheduling to approximate value functions and empirically outperforms discrete-time RL baselines; and Yildiz et al. (2021), which integrates neural ODEs with Bayesian inference to model uncertainty in state evolution and proposes a continuous-time actor–critic algorithm that mitigates challenges such as Q-function vanishing and poor discretization. In contrast, CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings. For example, Luviano and Yu (2017) solves the multi-agent pathfinding problem using fuzzy Q-learning, while Jiang et al. (2023) proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems. Beyond these examples, only a limited number of studies have addressed CT-MARL, highlighting the importance of our contributions.
@@ 2.2 Solving HJB Equations via PINNs
In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations (Crandall and Lions, 1983), which are the first-order nonlinear parabolic PDEs. However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD (Osher et al., 2004; Osher and Shu, 1991). Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth (Weinan et al., 2021). PINNs approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residuals (Han and Long, 2020; Han et al., 2018), PDE residuals (Bansal and Tomlin, 2021; Zhang et al., 2024a; b), and supervised data derived from numerical solvers (Nakamura-Zimmerer et al., 2021). Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments (Mukherjee and Liu, ). However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
@@ 3 Methodology
In this section, we present our Value Iteration via PINN (VIP) framework for solving CT-MARL problems. 1) We first formulate the problem setting as the cooperative multi-agent deterministic system in continuous-time. By adopting a CTDE paradigm, 2) we approximate the HJB-based value function using a PINN equipped with residual and anchor losses. To further enhance the stability and accuracy of both the value and its gradient, we incorporate additional VGI loss that iteratively refines both value and value-gradient accuracy along sampled trajectories. Building upon this critic, 3) we derive a continuous-time advantage function directly from the HJB residual to guide decentralized policy updates, which enables each agent to update its actor in a manner consistent with continuous-time optimality conditions.
@@ 3.1 Problem Formulation
In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework (Yildiz et al., 2021; Lee and Sutton, 2021), we formulate the continuous-time multi-agent problem as a tuple:
$\mathcal{M}=\bigl\langle\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,\{t_{k}\}_{k\geq 0},\rho\bigr\rangle.$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state space and $\mathcal{U}=\mathcal{U}_{1}\times\dots\times\mathcal{U}_{N}\subseteq\mathbb{R}^{m}$ represents the joint action space of $N$ agents. The global state and control input are represented by $x\in\mathcal{X}$ and $u\in\mathcal{U}$. Agent interactions occur over an infinite time horizon. The multi-agent system evolves according to time-invariant nonlinear dynamics defined by $\dot{x}=f(x,u)$, where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is the global dynamics function. We define $\pi:\mathcal{X}\rightarrow\mathcal{U}$ as the decentralized joint policy $\pi=(\pi_{1},\dots,\pi_{N})$. All agents share a global reward $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$. $\rho\in(0,1]$ is the discount factor or time-horizon scaling parameter. Unlike standard formulations that assume a fixed time step, we consider a strictly increasing sequence of decision times $\{t_{k}\}_{k\geq 0}$ with variable gaps $\tau_{k}=t_{k+1}-t_{k}>0$. In this paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f$ is Lipschitz continuous; $r$ is Lipschitz continuous and bounded.
@@ 3.2 HJB and Policy Learning
For CT-MARL problems, we build on HJB PDEs rather than discrete-time Bellman equations (Bellman, 1966), which are ill-suited to continuous-time settings. In this subsection, we explain how the HJB equations are leveraged to solve the CT-MARL problems. Specifically, we define a value network $V_{\theta}$ parametrized by a set of weights and biases $\theta$, and describe how the global value function $V_{\theta}$ is trained and how each agent’s policy $\pi_{\phi_{i}}$ is updated, such that the overall procedure serves as a continuous-time analogue of actor–critic policy iteration. The convergence of HJB-based PINNs for learning $V_{\theta}$ has been established in Meng et al. (2024), providing theoretical support for our approach.
3.2.1 Critic Learning with HJB
The optimal value function is fully represented by the HJB equation. This PDE encodes the fundamental relationship between reward, system dynamics, and value evolution, and serves as the theoretical backbone for our value learning method. The following lemma states this connection formally in the multi-agent cooperative setting.
@@ 3.2.1 Critic Learning with HJB
For all $x\in\mathcal{X}$, the value function $V(x)$ is the optimal solution to satisfy the following HJB PDEs:
$\rho V(x)+\nabla_{x}V(x)^{\!\top}\!f\bigl(x,u^{*}\bigr)+r\bigl(x,u^{*}\bigr)=0,$ (3)
where optimal control input $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,\nabla_{x}V(x))$. The Hamiltonian $\mathcal{H}$ is defined as $\mathcal{H}=\nabla_{x}V(x)^{\!\top}\!f\bigl(x,u\bigr)+r\bigl(x,u\bigr)$.
@@
To approximate differentiable value functions, we solve the HJB PDEs in Eq. 3. Since conventional numerical methods become intractable beyond six state dimensions (Bui et al., 2022), we instead employ PINNs, which approximate value functions by minimizing PDE residuals. Specifically, we define the HJB PDE residuals as
$\mathcal{R}_{\theta}(x_{t})=-\rho V_{\theta}(x_{t})+\nabla_{x_{t}}V_{\theta}(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}).$ (4)
and minimize the residual loss $\mathcal{L}_{\text{res}}=\bigl\|\mathcal{R}_{\theta}(x_{t})\bigr\|_{1}$ towards zero during model refinement.